OpenStudy (anonymous):

I don't understand integers. Why do -4 + -4 = -8 and -4 x -2 = 8 when both 4 + 4 = 8 and 4 x 2 = 8? Shouldn't both -4 + -4 and -4 x -2 be positive or negative?

OpenStudy (dumbcow):

ah the enigma of negative numbers here is a video that can do a better job than I https://www.khanacademy.org/math/pre-algebra/negatives-absolute-value-pre-alg/mult-div-negatives-pre-alg/v/why-a-negative-times-a-negative-is-a-positive

OpenStudy (anonymous):

P+P=P N+P=N N+N=N

OpenStudy (anonymous):

Thank you @dumbcow I'll watch it. @jque143 Is that the rule?

OpenStudy (anonymous):

Yeah its in the file